OpenStudy (anonymous):

Solve for the unknown: a+b+d+g+5+8+3+7=40 8+3+b+a=18 3+7+b+g=19 8+d+b+7=24 4 equations and 4 unknowns Please help me! :(

OpenStudy (anonymous):

we have to clean this up a bit first

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

\(a+b+d+g+5+8+3+7=40\) is the same as \(a+b+d+g = 17\)

OpenStudy (anonymous):

second one is the same as \[a+b = 7\]

OpenStudy (anonymous):

How did you get 7?

OpenStudy (anonymous):

\[8+3+b+a=18\] since \(8+3=11\) this is \[a+b+11=18\] subtract 11 and get \[a+b=7\]

OpenStudy (anonymous):

next one is \[b+g = 9\]

OpenStudy (anonymous):

and last one is \[b+d = 9\] that is what i meant by "clean it up" first

OpenStudy (anonymous):

I see. Thank you! What's the next step then?

OpenStudy (anonymous):

\[a+b+d+g=17\\a+b=7\\b+d=9\\b+g=9\] i think is what we have

OpenStudy (e.mccormick):

Ah, so the notes on the bottom of that other problem were not enough? What satellite73 did is the form I talked about. Now it should be very easy to get b. Once you have b, you have the rest.

OpenStudy (anonymous):

the last two equations tell you that \(d=g\) since if you subtract the one from the other you get \(d-g=0\) and so \(d=g\)

OpenStudy (anonymous):

then since \(a+b=7\) the first equation is \[7+d+g=17\] and so \[d+g=10\] now since \(d=g\) they both must be 5

OpenStudy (anonymous):

since \(d=5\) and \(b+d=9\) this tells you \(b=4\)

OpenStudy (anonymous):

and finally since you know \(b=4,d=g=5\) you can find \(a\) is that okay?

OpenStudy (anonymous):

YEEEEEEEEEEES! Thank you sooo much! :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

romantics?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=Rqnw5IfbZOU

OpenStudy (e.mccormick):

Now you can finish filling out that diagram. http://openstudy.com/study#/updates/51c1f48de4b0a49e8c8ff6fb

OpenStudy (anonymous):

wow, you don't need all this to solve that venn diagram!

OpenStudy (e.mccormick):

I was just continuing the way it was started. If there is an easier way, great. Add it. Would be good to see it.

OpenStudy (anonymous):

oh i see, i was looking at the first one, didn't see the second one sorry